Express the following as a product of disjoint cycle. Check them for being even or odd permutation and find the inverse of each of them in S7
1. (1 4 7) (2 6 5) (2 4 1) (5 6 7)
2. (7 1) (2 6 5) (2 4 1) (7 5 6)
1
Expert's answer
2020-11-17T05:43:11-0500

Let us express the permutations as a product of disjoint cycle using a rule for the product of cycles (from right to left): "1\to1\to2\to6\to 6, 6\to7\to7\to 7\to1" and so on. Therefore,

"(1 4 7) (2 6 5) (2 4 1) (5 6 7)=(16)(27)" is a product of disjoint cycles. This permutation consist of two transpositions, and thus is even. Since a transposition has order 2, the inverse of "(16)(27)" is "(27)(16)."


By analogy, "(7 1) (2 6 5) (2 4 1) (7 5 6)=(16)(247)" is a product of disjoint cycles. Since "(16)(247)=(16)(27)(24)", this permutation is odd. The inverse of "(7 1) (2 6 5) (2 4 1) (7 5 6)=(16)(27)(24)" is "(24)(27)(16)=(274)(16)."